# Prompts As Programs: A Structure-Aware Approach to Efficient Compile-Time Prompt Optimization

### v1

@@ 3.3 Specializations of Sammo
We note that Sammo is a rich framework that allows practitioners to mix-and-match search strategies with a large variety of mutation operators. The following methods are examples for special cases of Sammo:
  • APE – Automatic Prompt Engineering (Zhou et al., 2023). Here, InitializeCandidates generates a set of initial candidates from a small set of few-shot examples. Then, it uses a single mutation operator, Paraphrase with beam search to explore alternative candidates.
  • GrIPS – Gradient-free instruction search Prasad et al. (2023). This approach builds a syntax parse tree from the task instructions and then performs beam search with add, delete, swap and paraphrase mutation operations on the constituents.
4 Use Case: Instruction Tuning
Figure 3: Sammo consistently outperforms all other instruction tuning methods, independent of the backend LLM.
In theory, Sammo subsumes many existing methods that do instruction tuning where the task is to generate static instructions that tell the LLM how a task should be solved. To better align with previous work, we use eight zero-shot BigBench classification tasks Srivastava et al. (2023) that still have headroom to improve (i.e., the accuracy of the baseline prompt $\pi_{0}$ with GPT-3.5 is $<0.9$). We subsample training and test sets of size $n=100$ from the official splits.
We compare against APE (Zhou et al., 2023), Automatic Prompt Optimization (Pryzant et al., 2023) and GrIPS (Prasad et al., 2023). For the backend LLMs, we consider two open-source models, Mixtral 7x8B Jiang et al. (2024) and Llama-2 70B Touvron et al. (2023) and GPT-3.5 Brown et al. (2020). We do not include GPT-4 since it showed negligible headroom for improving instructions in pilot experiments (cf. also Section 5).
Results. As Figure 3 shows, Sammo is able to outperform all other baselines, independent of whether GPT-3.5, LLama-2 or Mixtral was used as a backend model. As a side note, the model baseline performance seems to be correlated with how much performance we gain through prompt tuning. Llama-2-70B sees largest relative performance gains (about 2x), Mixtral 7x8B moderate, and GPT-3.5 smallest gains (around 10% compared to the baseline instructions). However, the zero-shot tasks here are relatively simple, with more headroom existing for more complex tasks as the next use case shows.
5 Use Case: Optimizing Retrieval Augmentation
@@
Figure 5: There is only weak correlation between how well candidates do during enumerative search between LLMs.
Towards a more realistic application of prompt optimization, we consider improving retrieval-augmented semantic parsing. The overall task is to translate a natural user query into a domain-specific language (DSL) construct. Our goal is to show how application of Sammo can yield substantial gains in a realistic few-shot scenario. To stay within our overall setting of limited labeled data availability, we create a training/few-shot split and test split with ($n=500$ and $n=100$ each). To measure candidate performance, we subsample 100 examples from the training split as an evaluation set. To the best of our knowledge, there are no existing baselines for this scenario.
Following Bogin et al. (2023), we use three different datasets/domains: GeoQuery (Zelle and Mooney, 1996), SMCalFlow (Andreas et al., 2020) and Overnight (Wang et al., 2015). We use the same i.i.d. splits, DSLs and starting prompt format as Bogin et al. (2023). We used Sammo with enumerative search to optimize exact match accuracy over a search space of size 24, trying out different data formats, number of few shot examples and DSL specifications. For details, see Appendix A.5. We consider all backend LLMs as before, plus GPT-4.
Results. As Figure 4 shows, despite its conceptual simplicity, optimizing retrieval-augmented prompts with Sammo yields substantial gains across most datasets and backend LLMs. We note that as before, relative gains decrease with increasing model strength. Llama-2 sees an average improvement of 133% and Mixtral of 44%. However, even GPT-4 can benefit from changes explored by Sammo with a average relative gain of 30%.
@@
6 Use Case: Prompt Compression
In these experiments, our goal is to optimize the weighted costs of a metaprompt subject its accuracy not dropping below a certain threshold $\tau_{acc}$ and its parse error rate being below 10%.
The weighted costs are given as
$\text{cost}(\pi,D)=S_{\text{in}}(\pi,D)+2\cdot S_{\text{out}}(\pi,D).$ (4)
where we weight the number of output tokens double output tokens $S_{\text{out}}$ compared to the input tokens $S_{\text{in}}$ to reflect current billing schemes for publicly available LLMs, but the loss function is easily configurable upfront. Assuming a baseline prompt $\pi_{0}$ that has a reasonable level of accuracy, we set the threshold $\tau_{\text{acc}}$ such that it is above the baseline prompt performance with $\epsilon=0.02$.
We sampled ten classification tasks with longer instructions (1000 characters or more) from the Super-NaturalInstructions benchmark (Wang et al., 2022).
@@ 6 Use Case: Prompt Compression
Baselines. To make the compression task realistic and start with a competitive prompt, we batch input examples with the newline delimited format of Cheng et al. (2023). Based on pilot experiments, we chose input batch sizes $b$ such that performance between no batching and input batching was within $\epsilon$. This resulted in $b=10$ for GPT-4, $b=5$ for Mixtral and GPT-3.5, and $b=1$ for Llama-2. We compare Sammo against four other compile-time prompt compression techniques:
  • Baseline Prompt: Our baseline prompt $\pi_{0}$ uses the official instructions provided with a task followed by $k=5$ in-context examples.
  • APE: Automatic Prompt Engineering (Zhou et al., 2023): Instead of optimizing for accuracy, we used the same objective as for Sammo (Eq. 4)
  • STDC: Syntax-guided Task Definition Compression (Yin et al., 2023a) runs a sequential search to prune syntax tree constituents.
  • Stopwords: We search over two different stop word lists as well as whether to apply them to examples, to the task definition, or both (implemented as RemoveStopwords operators from Table 1).
  • GPT-4 Rewrite: Using the templates from Li et al. (2023), we try out ten different instructions to compress the task definition and the in-context examples.
  • Sammo: We use beam search as the backbone of Algorithm 1 and initialize $\Pi_{C}$ with four variants for $\theta_{\text{format}}$ of the baseline prompt $\pi_{0}$. We use all mutation operators listed in Table 1 as possible operations, and choose mutators uniformly at random during the search.
Results. Figure 6 show the the average performance across the ten tasks for all backbone LLMs. We show the final weighted costs on the test set (left y-axis), as well as the difference of performance relative to the baseline prompt which should ideally not exceed $\epsilon=0.02$ (right y-axis).
For all back-end models, Sammo achieves substantial compression rates, reducing the costs by over 40% while maintaining the accuracy of the baseline prompt. The STDC and Stopwords baselines achieve some compression, but the compression rates seen are only moderate, most likely because their mutation operations are limited. APE and GPT-4 Rewrite manage to compress prompts to a larger degree, but can find prompts that do not generalize well to the test set and experience large drops in accuracy.
@@
For each mutation operation chosen by Sammo during the search process, we recorded whether it resulted in an improvement of the objective (Eq. 4). This gives us a rough idea of how much individual operators contribute to the success of the search. Figure 7 shows the fraction of times each operator resulted in an improvement when it was chosen. From that, we can see that how successful a mutator is depends on the backend LLM, but rewriting operations and dropping in-context examples were the most useful structural mutators compressing the prompt. GPT-4’s performance was more robust to lowering the number of in-context examples, and also to dropping the introduction than other backend LLMs.
@@ 7 Related Work
In prompt compression, one main axis of distinction is what model access methods assume. Assuming full model access, prompt tuning learns a mapping function that translates an initial prompt to either soft or hard tokens for efficiency (e.g., Lester et al. (2021)). For example, Wingate et al. (2022) uses a distillation objective to steer text generation with compressed instructions and Mu et al. (2023) use meta-learning to compresses prompts into “gist” tokens.
@@
In prompt optmization, the main focus in on improving the accuracy of prompts. Past work has typically focused on simpler (e.g. single task, non-batched) prompts with less structure. Again, there are a variety of methods that assume full model access  (Lester et al., 2021; Qin and Eisner, 2021) which we will not discuss further. Working with continuous prompt representations using a smaller surrogate model, InstructZero Chen et al. (2023) optimizes prompts locally via Bayesian optimization and uses calls to the LLM API as feedback. The main limitation here is similar to token-level methods; it is unclear how to apply them to structure-rich prompts. On the discrete optimization side, Automatic Prompt Engineer (APE) Zhou et al. (2023) generates instruction candidates from a few input-output pairs, and then uses beam search over paraphrased candidates. We use a modified version with the same objective as Sammo in our experiments. Targeting mainly accuracy and not compression, GrIPS Prasad et al. (2023) uses beam search with edit, add and delete operations on the syntax tree after parsing the instructions. Similarily, Automatic Prompt Optimization (APO) Pryzant et al. (2023) re-writes instructions by generating explanations for errors, changing the prompt to reflect explanations, and then generating more prompt candidates by paraphrasing.
@@
All of our experiments have been carried out with datasets in English; performances for lower-resource language are likely to be lower. While Sammo is generally efficient, tasks need to have a certain level of downstream usage in order to compensate for the upfront costs of optimization. Finally, Sammo adopts a supervised learning scenario where labels are required; we plan to address more unsupervised tasks in the future.
8 Conclusion
In this paper, we introduced a new framework, Structure-aware Multi-Objective Metaprompt Optimization (Sammo) to enable efficient optimization of metaprompt programs during compile-time.
Sammo represents metaprompts as dynamic function graphs, and employs a set of mutation operators to alter the structure and content of metaprompts. This approach notably outperforms and generalizes existing methods of prompt optimization and compression, as demonstrated through several use cases tasks in our experimental evaluation.
@@ A.5 RAG optimization
A.5.1 Baseline

### v2

@@ 4.3 Specializations of Sammo
We note that Sammo is a rich framework that allows practitioners to mix-and-match search strategies with a large variety of mutation operators. The following methods are examples of how known methods can be implemented in Sammo:
APE
– Automatic Prompt Engineering (Zhou et al., 2023). Here, InitializeCandidates generates a set of initial candidates from a small set of few-shot examples. Then, it uses a single mutation operator, Paraphrase with beam search to explore alternative candidates.
GrIPS
– Gradient-free instruction search Prasad et al. (2023). This approach builds a syntax parse tree from the task instructions and then searches with Add, Delete, Swap and Paraphrase mutators on the constituents.
5 Experiments
We demonstrate the flexibility and effectiveness of Sammo’s SPP search approach across three different scenarios, comparing to the most appropriate baselines in each scenario. In line with our assumption of limited data availability (Section 3), we sample $n=100$ examples for training and test sets each unless noted otherwise. For the backend LLMs, we consider two open-source models, Mixtral 7x8B Jiang et al. (2024) and Llama-2 70B Touvron et al. (2023); as well as two closed-source models, GPT-3.5 and GPT-4 Brown et al. (2020). See Appendix A.1 for model details. Unless noted otherwise, we use Sammo with beam search and a search budget (fixed for all baselines) of $B=48$ candidate evaluations.
5.1 Instruction Tuning
Figure 3: Sammo consistently outperforms all other instruction tuning methods, for all of the backend LLMs.
To better align with previous work on prompt tuning, we ran our experiments on BigBench zero-shot classification tasks Srivastava et al. (2023). We sampled eight tasks that still had headroom to improve, i.e., tasks where the baseline prompt $\pi_{0}$ with GPT-3.5 had an accuracy of $<0.9$.
We compare against APE (Zhou et al., 2023), APO (Automatic Prompt Optimization, Pryzant et al., 2023), DSpy COPRO Khattab et al. (2023) and GrIPS (Prasad et al., 2023). We do not include GPT-4 since it showed negligible headroom for improving instructions in these simple prompt programs in pilot experiments (cf. also Section 5.2).
Results. As Figure 3 shows, Sammo is able to outperform all other baselines, independent of whether GPT-3.5, LLama-2 or Mixtral was used as a backend model. DSpy COPRO Khattab et al. (2023) performed even worse than the baseline prompt since DSPy’s prompting often caused the model to not adhere to the output format (see Appendix A.2.1 for an example). As a side note, the model baseline performance seems to be correlated with how much performance we gain through prompt tuning. Llama-2-70B sees largest relative performance gains (about 2x), Mixtral 7x8B moderate, and GPT-3.5 smallest gains (around 10%) compared to the baseline instructions.
5.2 Optimizing Retrieval Augmentation
@@
Figure 5: There is only weak correlation between how well enumerative search candidates do across LLMs.
Towards a more realistic application of prompt optimization, we consider improving retrieval-augmented semantic parsing. The overall task is to translate a natural user query into a domain-specific language (DSL) construct. Here, we sample 500 examples for retrieval, and use a subset of 100 examples from those to evaluate training performance. We compare to DSpy MIPRO Khattab et al. (2023), a method that optimizes few-shot examples as well the instructions. Since evaluation is expensive in this scenario due to longer prompts, we limit the budget to $B=24$ candidate evaluations.
Following Bogin et al. (2023), we use three different datasets/domains: GeoQuery (Zelle and Mooney, 1996), SMCalFlow (Andreas et al., 2020) and Overnight (Wang et al., 2015). We use the same i.i.d. splits, DSLs and starting prompt format as Bogin et al. (2023). We used Sammo with enumerative search to optimize the data formats, number of few shot examples and DSL specifications. For details, see Appendix A.5.
Results. As Figure 4 shows, despite its conceptual simplicity, optimizing retrieval-augmented prompts with Sammo yields substantial gains across most datasets and backend LLMs. We note that as before, relative gains decrease with increasing model strength. Llama-2 sees an average improvement of 133% and Mixtral of 44%. However, even GPT-4 can benefit from changes explored by Sammo with a average relative gain of 30%. DSpy MIPRO does worse than Sammo in all but one setting. Besides outputting answers in the wrong format, MIPRO also suffers from overfitting as training accuracies can reach 100% (see Appendix A.5.1).
@@
5.3 Prompt Compression
In these experiments, our goal is to optimize the weighted costs of a prompt program while maintaining its accuracy.
The primary objective is the weighted sum of the number of input tokens with weight one and the number of output tokens with weight two to reflect current billing schemes of popular LLM providers. Assuming a baseline prompt $\pi_{0}$ that has a reasonable level of accuracy, we set a threshold $\tau_{\text{acc}}$ such that the performance of the compressed prompt is required to be above the baseline prompt performance with margin $\epsilon=0.02$. We sampled ten classification tasks with longer instructions (1000 characters or more) from the Super-NaturalInstructions benchmark (Wang et al., 2022).
@@ 5.3 Prompt Compression
Baselines. To make the compression task realistic and start with a competitive prompt, we batch input examples with the newline delimited format of Cheng et al. (2023). Based on pilot experiments, we chose input batch sizes $b$ such that performance between no batching and input batching was within $\epsilon$. This resulted in $b=10$ for GPT-4, $b=5$ for Mixtral and GPT-3.5, and $b=1$ for Llama-2. We compare Sammo against four other compile-time prompt compression techniques. Our baseline prompt $\pi_{0}$ uses the official instructions provided with a task followed by $k=5$ in-context examples. For Sammo, we use all mutation operators listed in Table 1 as possible operations, and choose mutators uniformly at random during the search. We also compare against APE, with the cost-aware objective. APO is not applicable since it optimizes only for accuracy. New baselines that were added include:
STDC
– Syntax-guided Task Definition Compression (Yin et al., 2023a) runs a sequential search to prune syntax tree constituents.
Stopwords
– This is using Sammo limited to the RemoveStopwords operators from Table 1.
GPT-4 Rewrite
– Using the templates from Li et al. (2023), we try out ten different templates to shorten the instructions.
Results. Figure 6 show the the average performance across the ten tasks for all backbone LLMs. We show the final weighted costs on the test set (left y-axis), as well as the difference of performance relative to the baseline prompt which should ideally not exceed $\epsilon=0.02$ (right y-axis). For all back-end models, Sammo achieves substantial compression rates, reducing the costs by over 40% while maintaining the accuracy of the baseline prompt. The STDC and Stopwords baselines achieve some compression, but the compression rates seen are only moderate, most likely because their mutation operations are limited. APE and GPT-4 Rewrite manage to compress prompts to a larger degree, but can result in prompts that do not generalize well to the test set and experience large drops in accuracy.
@@
For each mutation operation chosen by Sammo during the search process, we recorded whether it resulted in an improvement of the weighted costs. This gives us a rough idea of how much individual operators contribute to the success of the search. Figure 7 shows the fraction of times each operator resulted in an improvement when it was chosen. From that, we can see that how successful a mutator is depends on the backend LLM, but rewriting operations and dropping in-context examples were the most useful structural mutators compressing the prompt. GPT-4’s performance was more robust to lowering the number of in-context examples, and also to dropping the introduction, than other backend LLMs.
@@ 6 Related Work
In prompt compression, one main axis of distinction is what model access methods assume. Assuming full model access, compression via prompt tuning learns a mapping function that translates an initial prompt to either soft or hard tokens for efficiency (e.g., Lester et al. (2021)). For example, Wingate et al. (2022) uses a distillation objective to steer text generation with compressed instructions and Mu et al. (2023) use meta-learning to compresses prompts into “gist” tokens.
@@
In prompt optimization, the main focus in on improving the accuracy of prompts. Past work has typically focused on simpler (e.g. single task, non-batched) prompts with less structure. Again, there are a variety of methods that assume full model access  (Lester et al., 2021; Qin and Eisner, 2021) which we will not discuss further. Working with continuous prompt representations using a smaller surrogate model, InstructZero Chen et al. (2023) optimizes prompts locally via Bayesian optimization and uses calls to the LLM API as feedback. The main limitation here is similar to token-level methods; it is unclear how to apply them to structure-rich prompts. On the discrete optimization side, Automatic Prompt Engineer (APE) Zhou et al. (2023) generates instruction candidates from a few input-output pairs, and then uses beam search over paraphrased candidates. We use a modified version with the same objective as Sammo in our experiments. Targeting mainly accuracy and not compression, GrIPS Prasad et al. (2023) uses beam search with edit, add and delete operations on the syntax tree after parsing the instructions. Similarily, Automatic Prompt Optimization (APO) Pryzant et al. (2023) re-writes instructions by generating explanations for errors, changing the prompt to reflect explanations, and then generating more prompt candidates by paraphrasing.
7 Conclusion
In this paper, we introduced Sammo, a framework for efficient compile-time optimization of prompt programs. The key innovation of Sammo is to represent prompts as symbolic prompt programs. Symbolic prompt programs enable Sammo to search through the space of possible programs in an efficient manner. This approach notably outperforms and generalizes existing methods of prompt optimization and compression, as demonstrated through several use cases tasks in our experimental evaluation.
Sammo is made available publicly as an open-source project. It is our hope that similar to how compilers accelerated and robustified software programming, prompt program compilers like Sammo will accelerate prompt development, model evaluation and prompt migration between different architectures.
@@ 7 Conclusion
The optimization costs as well as the resulting prompt program performance could be sensitive to search hyperparameter choices. Given the high cost of running just one experiment, we could not afford to run a full hyperparameter search, but did our best to choose settings based on recommendations by the authors or previous similar work and ensured that all algorithms had the same budget.
Due to Sammo high-level operators, we did not observe substantial drops in performance between training and test sets. However, methods that mostly optimize in-context examples like DSPy showed a large risk of overfitting (cf. Section 5.2). We conclude that more research is needed to understand when and how overfitting occurs prompt optimization.
All of our experiments have been carried out with datasets in English; performances for lower-resource language are likely to be lower. While Sammo is generally efficient, tasks need to have a certain level of downstream usage in order to compensate for the upfront costs of optimization. Future research could also combine run-time optimization with compile-time optimization to see if additional gains can be achieved. Finally, Sammo adopts a supervised learning scenario where labels are required; we plan to address more unsupervised tasks in the future.
@@ A.2 Instruction Tuning
A.2.1 DSPy Example Output
DSpy often caused the model to not adhere to output instructions.
  Input      Does Speaker 2’s answer mean yes or no?  ---  Follow the following format.  Question: \${question}  Answer: \${answer}  ---  Question: Speaker 1: ’Should we repay you?’ Speaker 2: ’There is no charge for awesomeness, or attractiveness.’  Answer:
  Output       No  Question: Speaker 1: ’Should we repay you?’ Speaker 2: ’There is no charge for awesomeness, or attractiveness.’  Answer: No  Question: Speaker 1: ’Should we repay you?’ Speaker 2: ’There is no charge for awesomeness, or attractiveness.’  Answer: No  Question: Speaker 1: ’Should we repay you?’ Speaker 2: ’There is no charge for awesomeness, or attractiveness.’  Answer: No  Question: Speaker 1: ’Should we repay you?’ Speaker 2: ’There is no charge
@@ A.5 RAG optimization
A.5.1 Overfitting in DSpy MIPRO
See Figure 8.
Figure 8: Training accuracies on RAG task
A.5.2 Baseline